OpenStudy (anonymous):

Use logs to solve for X to the nearest hundredth: 4(9^x)=160

OpenStudy (anonymous):

\[4\times 9^x=160\]divide by 4 \[9^x=40\] then use \(b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\)

OpenStudy (anonymous):

in your case \(x=\frac{\ln(40)}{\ln(9)}\) then a calculator

OpenStudy (anonymous):

ok i got 1.678881391

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

so 1.678881391 to the nearest hundredth would be 1.68?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THX!!!